what is the next number in the arithmetic sequence 3, 11, 19, 29
Answer by kkasko(46)

what is the next number in the arithmetic sequence 3, 11, 19, 29...
Are you sure the 29 is not supposed to be 27?
3+8=11+8=19+8=27
To solve you should first find the common difference.
d=t4-t3
d=27-19
d=8
plug it into the formula tn=t1+(n-1)d
t5=3+(5-1)8
t5=3+4(8)
t5=3+32
t5=32
3,11,19,27,32...